OpenStudy (studybird123):

What is the exact area of a circular plate with a radius of 15 cm? A. 15π cm2 B. 30π cm2 C. 225π cm2 D. 900π cm2

OpenStudy (anonymous):

use 2times pie r

OpenStudy (anonymous):

multiply 2 times 15

OpenStudy (anonymous):

\[area=\pi r^2\]

OpenStudy (studybird123):

Oh thank you! Sorry I wasn't here!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

area of circular plate = pi r^2 here r = 15 so r^2 =15^2 = 225 so the area of circular plate = 225 pi. cm^2 so the option C